OpenStudy (anonymous):

Evaluate –6(x – 2y) when x = –3 and y = –5

OpenStudy (nikkibxox):

Plug it in? Do youknow how

OpenStudy (anonymous):

–6(x – 2y) just put the value of x and y to this above equation

OpenStudy (anonymous):

–6(x – 2y) = -6 (-3 -2 (-5)) =-6(-3+10) =-6*7 =-42 @katrina19123